# Fine-Tuning Code Language Models to Detect Cross-Language Bugs

### v1

@@ 5.1. Interpretation of Study Results
RQ1: CodeLM performance after fine-tuning: This study reveals significant performance differences among CodeLMs of various sizes in CLB detection tasks. Although slight decreases were observed in some evaluation metrics after fine-tuning, overall, nearly all CodeLMs achieved varying degrees of performance improvement. This observation suggests that the CLB dataset we constructed effectively enhances CodeLMs’ modeling capabilities in handling CLB detection challenges.
@@
Notably, the size of model parameters exhibits a considerable impact on the performance of CodeLMs fine-tuned for CLB detection. As shown in Tables 2 and 4, CodeLMs with a relatively smaller number of parameters (no more than 220M) generally achieved better results, while large CodeLMs, such as Qwen2.5_Coder-7B and CodeLlama-7b-hf, demonstrated only moderate performance. This finding indicates that solely increasing the model size does not necessarily guarantee superior performance in CLB detection tasks. We speculate that small CodeLMs, due to their more streamlined architecture and more focused training, are more amenable to fine-tuning and adaptation for targeted tasks such as CLB detection (Xu et al., 2025; Hsieh et al., 2023; Ahouzi et al., 2024).
Furthermore, UniXcoder-base is ranked first across all evaluation metrics, highlighting that the depth of a CodeLM’s understanding of code structure and semantics during pre-training plays a crucial role in cross-language tasks. Therefore, the pre-training strategy and architectural design of a CodeLM could be more critical factors than merely scaling up the number of parameters, offering valuable insights for the future design of efficient CodeLMs.
RQ2: Performance of CodeLMs fine-tuned on single-language bug datasets for CLB detection: Although CodeLMs fine-tuned on single-language bug datasets such as CVEFixes and CodeNet perform well in the single-language bug detection task, they exhibit suboptimal performance in CLB detection, generally showing lower F1 scores and AUC values. This phenomenon indicates a fundamental difference between single-language bugs and CLBs.
CLBs often arise from inconsistencies or misuse in the interaction mechanisms between two or more PLs, and this CLB-specific complexity cannot be reflected in single-language bug datasets. As a result, CodeLMs trained solely on single-language bug datasets experience struggle to capture the potential issues inherent in cross-language interactions.
The findings of this study further emphasize the importance of constructing dedicated CLB datasets. They also highlight the need to develop more targeted training methods and model adaptation strategies specifically for CLB detection tasks. Future model training efforts should pay particular attention to cross-language factors, such as language boundary issues and differences in interaction patterns.
@@
For CodeLMs that showed notable performance improvements, code comments provided additional context and explanations that contributed to a more accurate understanding of the code, suggesting that these CodeLMs possess relatively limited inherent code understanding capabilities. Therefore, in practical applications, whether to retain code comments should be flexibly decided based on the specific task requirements. It is also advisable to consider adopting more flexible token budget strategies, such as moderately increasing the token sequence length limit or prioritizing key code content. This finding further indicates that current CodeLMs still have considerable potential for improvement in code understanding tasks.
@@ 5.2. Implications
Based on the study results and their analysis, we obtained a number of implications for practitioners and researchers, respectively.
@@ 5.2.1. Implications for practitioners
(1) Emphasizing the specific characteristics of CLBs. This study shows that CLBs are not only considerable in quantity but also tend to be more obscure and harder to detect compared to single-language bugs, due to the involvement of interaction mechanisms across different languages. Therefore, in practical software development, it is recommended to consider potential risks related to cross-language invocations during the design phase, such as data type incompatibility, inconsistent exception handling, and differences in resource management. Moreover, during testing and code review phases, it is advisable to introduce specialized checklists and test cases targeting cross-language interaction scenarios to improve bug detection performance. (2) Paying attention to limitations of CodeLMs in CLB detection. Although CodeLMs have made significant progress in tasks such as code understanding and generation, their performance on CLB detection remains limited. Practitioners should not regard the outputs of CodeLMs as final conclusions when detecting CLBs. Instead, a multifaceted and multi-perspective detection process should be adopted by combining static analysis tools, rule-based detection methods, and manual review to comprehensively assess the possibility of CLBs. Over-reliance on the output of a single CodeLM should be avoided to enhance overall reliability of CLB detection. (3) Leveraging code comments appropriately to avoid information loss. In this study, the results show that including code comments generally improves detection performance. However, it may also cause information loss due to input token sequence length exceeding the model’s limit. Therefore, in practical CLB detection application, it is recommended to dynamically adjust the token budget based on code length or design a comment selection mechanism that retains only those comments closely related to functionality and critical for understanding code. This way aims to maximize the positive impact of comments on CLB detection performance while mitigating potential negative effects. (4) Selecting an appropriate CodeLM based on resource availability and projects under detection. The results indicate that large CodeLMs are not necessarily superior to small ones in CLB detection tasks. Depending on resource availability and the specific projects under detection, small models with fewer parameters and task-specific training can be prioritized. Such CodeLMs offer lower deployment costs while maintaining competitive performance, making them well-suited for scenarios with limited resources or strict latency requirements.
@@ 5.2.2. Implications for researchers
(1) Expanding the coverage of CodeLMs and tasks. The current study evaluates a limited number of CodeLMs, with different model architectures and parameter sizes. Future work should systematically expand the research scope to include a broader range of CodeLM architectures and large CodeLMs. Additionally, the transferability and robustness of CodeLMs across diverse cross-language scenarios should be explored to establish more comprehensive and accurate evaluation standards. (2) Building larger and more diverse CLB dataset. This study provides evidence of the importance of dataset quality and fine-tuning strategies. However, existing CLB datasets are still limited in terms of size, language diversity, and interaction patterns. Future work should focus on constructing larger and more comprehensive benchmark datasets that cover a wider range of PL combinations and cross-language interaction modes, in order to systematically evaluate the detection capabilities of CodeLMs of different scales and structures. (3) Investigating the fundamental characteristics and classification of CLBs. Currently, the understanding of CLBs remains relatively shallow, and there is a lack of systematic classification standards regarding the various types of CLBs. Future research could aim to establish a structured classification system for CLBs based on their sources (e.g., memory management, exception handling, data format conversion) and manifestations (e.g., data loss, type errors, resource leaks). Such a classification system would not only facilitate more targeted CodeLM training but also support the development of finer-grained CLB detection strategies. (4) Exploring fine-tuning strategies and data augmentation methods for CLB detection. While it is shown that increasing the volume of training data has improved CodeLM performance, efficiently leveraging limited data remains a significant challenge. Future directions include exploring self-supervised learning, domain adaptation, and cross-language contrastive learning to enhance model performance under low-resource conditions. Additionally, designing smarter data augmentation methods, such as automatically generating diverse CLB cases, could further enrich the training dataset and ultimately improve the CodeLMs’ performance in detecting CLBs.
@@ 6.2. External Validity
To answer RQ1 and RQ4, we fine-tuned CodeLMs on datasets with and without code comments, respectively, to investigate their performance in CLB detection tasks and to analyze the impact of code comments on CodeLMs effectiveness. A major threat to external validity in this setting lies in the limited parameter sizes of the CodeLMs used. In this study, we fine-tuned only 13 relatively small CodeLMs (no larger than 7B). As a result, the findings may not generalize to the performance of large models in CLB detection. Furthermore, due to computational constraints, we employed LoRA, a parameter-efficient fine-tuning (PEFT) technique, for CodeLM variants with more than 770M parameters. This may affect the generalizability of our study findings.
@@ 6.3. Internal Validity
A key internal threat lies in the choice of hyper-parameters during the fine-tuning process. The hyper-parameters of these 13 CodeLMs during the fine-tuning process play a crucial role in the final performance on the CLB detection task. Changes in hyper-parameters can lead to different results, potentially affecting the reliability of our experimental findings. To mitigate this threat, we conducted multiple test experiments to determine the optimal hyper-parameters.
Another threat stems from the different fine-tuning strategies used for CodeLMs of different sizes: small CodeLMs were fine-tuned using full training, while large CodeLMs were fine-tuned using PEFT. This inconsistency may affect the fairness of model comparisons. To alleviate this threat, we explicitly reported the fine-tuning strategy used for each model in our results, and we highlighted these differences in our comparative analysis. When interpreting performance differences, we took the fine-tuning method into account to avoid drawing misleading conclusions based on absolute scores.
@@ 6.4. Reliability
To further ensure the reliability and reproducibility of our experiments, we publicly shared the data collection standards and the final CLB dataset (Li et al., 2025). Detailed descriptions were provided for each step of the data collection process (see Section 3.3), including repository selection criteria (e.g., repositories with more than 500 stars, cross-language collaboration, and relevant bug issues) and the procedure of data processing. These steps ensured the standardization of data collection, and all relevant parameters (such as the repository star threshold and the minimum proportion of code in each programming language) were clearly specified and justified in the methodology.
In the dataset partitioning phase, we used a fixed random seed to ensure consistent splitting of the training, validation, and test sets, thereby eliminating randomness that could affect the results. Additionally, we made our experimental scripts publicly available(Li et al., 2025) to guarantee the reproducibility of our experiments. Thus, with these measures, the threats to reliability are alleviated.
@@ 7. Conclusions and Future Work
In this work, we designed CLCFinder, a tool for identifying cross-language code, and constructed a real-world dataset for CLB detection from GitHub. Based on this dataset, we conducted experiments to evaluate the performance of widely-used CodeLMs in the CLB detection task. The results show significant performance variation across different CodeLMs, with small CodeLMs (the parameter size is 220M or less) often outperforming large CodeLMs. To further validate the differences between our CLB dataset and single-language bug datasets, we fine-tuned CodeLMs on SOTA single-language bug detection datasets and applied these fine-tuned CodeLMs to CLB detection. The results demonstrate significant differences in feature characteristics between single-language bug and CLB datasets, validating the effectiveness and necessity of our CLB dataset. However, although the mainstream CodeLMs have demonstrated preliminary effectiveness in CLB detection, they are still primarily used in research settings and are not yet ready for real-world applications. We then explored factors affecting the performance of CodeLMs in CLB detection. The results show that increasing dataset size improves CodeLM performance, while the impact of token sequence length varies across CodeLMs. Some CodeLMs perform better with a shorter token sequence length, while others perform better with a longer token sequence length. Finally, we investigated the impact of code comments on model performance in CLB detection. Experimental results indicate that code comments affect CodeLMs differently - some CodeLMs perform better with commented code, some CodeLMs show no significant performance change, and others perform worse.
@@
For future work, we will advance research on CLB detection along several key directions. First, we plan to expand our dataset beyond Python and Java to include more PLs. This will enable more comprehensive CLB studies and improve the generalizability of detection models. Additionally, since our dataset includes bugs related to high-severity security vulnerabilities, we aim to explore cross-language security vulnerability detection in future research. Second, while this study focused on relatively small CodeLMs (no larger than 7B), we plan to investigate the performance of large CodeLMs in CLB detection. We also intend to conduct more in-depth analysis to better understand the factors that influence CLB detection performance, such as model architecture and language-specific features. Finally, we will examine the effect of longer input sequences and larger datasets on model performance. In this study, input token sequence length was limited to 512, which may have restricted the CodeLM’s ability to capture long-range dependencies. We plan to expand our dataset in order to study the impact of dataset size on CLB detection.
@@
The replication package of this work has been made available at (Li et al., 2025).

### v2

@@ 5.1. Interpretation of Study Results
RQ1: CodeLM performance after fine-tuning: This study reveals significant performance differences among CodeLMs of various sizes in CLB detection tasks. Overall, the evaluated CodeLMs exhibited varying levels of performance after task-specific adaptation, with some models achieving competitive results while others remained relatively modest. This observation suggests that the CLB dataset we constructed effectively enhances CodeLMs’ modeling capabilities in handling CLB detection challenges.
@@
Notably, the size of model parameters exhibits a considerable impact on the performance of CodeLMs fine-tuned for CLB detection. As shown in Tables 2 and 8, under our experimental setup, CodeLMs with a relatively smaller number of parameters (no more than 220M) generally achieved better results, while large CodeLMs, such as Qwen2.5_Coder-7B and CodeLlama-7b-hf, demonstrated only moderate performance. It is important to note that this observation is made under specific conditions, where smaller models are fully fine-tuned, several larger models are adapted using LoRA, and all models are evaluated with a maximum input length of 512 tokens. Therefore, within this setting, solely increasing the model size does not necessarily guarantee superior performance in CLB detection tasks. We speculate that under these constraints, small CodeLMs, due to their more streamlined architecture and more focused training, may be more amenable to fine-tuning and adaptation for targeted tasks such as CLB detection (Xu et al., 2025; Hsieh et al., 2023; Ahouzi et al., 2024).
In addition, the inclusion of GPT-4o-mini as an inference-only reference baseline provides complementary insights into the role of task-specific adaptation. As an inference-only general LLM, GPT-4o-mini is consistently outperformed by fine-tuned CodeLMs across all evaluation metrics. This result indicates that, under our experimental setup, prompt-based reasoning alone, without parameter adaptation, may be insufficient to fully capture the complex cross-language interaction patterns required for accurate CLB detection.
Furthermore, UniXcoder-base is ranked first across all evaluation metrics, highlighting that the depth of a CodeLM’s understanding of code structure and semantics during pre-training plays a crucial role in cross-language tasks. The contrast between fine-tuned CodeLMs and the inference-only GPT-4o-mini baseline further suggests that effective CLB detection in this setting may benefit more from task-specialized representation learning than from prompt-based general-purpose reasoning capabilities alone. However, we interpret this comparison cautiously. As a relatively compact and earlier-generation model, GPT-4o-mini serves as a helpful reference point for lightweight inference but does not represent the capabilities of state-of-the-art frontier LLMs (e.g., GPT-4 or Claude 3.5 Sonnet). Consequently, this baseline illustrates the performance gap between non-adapted general models and fine-tuned domain models of similar scales, but it does not reflect the detection capabilities of current state-of-the-art models on the CLB dataset. These findings suggest that, in our setting, the pre-training strategy and architectural design of a CodeLM may play a more important role than merely scaling up the number of parameters, offering valuable insights for the future design of efficient CodeLMs.
RQ2: Impact of Different Fine-Tuning Datasets on CLB Detection Performance: The results of this RQ suggest that fine-tuning CodeLMs on single-language bug datasets provides limited benefits for CLB detection. While such datasets are effective for learning language specific bug patterns, they do not adequately support capturing the characteristics of cross-language interactions required for CLB detection.
This limitation can be attributed to the fundamental differences between single-language bugs and CLBs. Single-language bug datasets mainly capture defects arising from syntax, semantics, or control flow within an individual PL. In contrast, CLBs typically originate from mismatches and incorrect assumptions at language boundaries, such as inconsistent data representations, interface misuse, or incompatible interaction protocols between multiple PLs. These cross-language characteristics are inherently absent from single-language training data, making it difficult for CodeLMs to learn transferable representations for CLB detection.
The findings highlight that effective CLB detection depends not only on CodeLM capacity but also on the relevance of the training data. Fine-tuning on CLB dataset allows CodeLMs to directly observe and learn interaction patterns across PLs, which is essential for detecting CLBs. Therefore, relying solely on single-language bug datasets is insufficient for addressing the unique challenges posed by cross-language software systems.
To further examine whether CodeLMs fine-tuned on the CLB dataset can transfer to single-language bug detection tasks, we conduct experiments on two single-language datasets, CVEfixes and CodeNet, together with our CLB dataset. Specifically, for each single-language dataset, we fine-tune two CodeLMs on the CLB dataset and on the target single-language dataset, respectively. We then evaluate all fine-tuned CodeLMs on the test set of the corresponding single-language datasets. This experimental setting is consistent with that used in RQ2.
Table 12. Single-language Bugs Detection Performance of CodeLMs Fine-Tuned on CVEfixes and the CLB Dataset
CodeLM Fine-tuned in CLB Fine-tuned in CVEfixes
Acc Prec Rec F1 AUC Acc Prec Rec F1 AUC
CodeBERT-base 0.4923 0.4908 0.4065 0.4447 0.4776 0.6352 0.6000 0.8112 0.6898 0.6867
UniXcoder-base 0.5051 0.5027 0.9388 0.6548 0.4829 0.6480 0.6033 0.8639 0.7105 0.7021
Table 13. Single-language Bugs Detection Performance of CodeLMs Fine-Tuned on CodeNet and the CLB Dataset
CodeLM Fine-tuned in CLB Fine-tuned in CodeNet
Acc Prec Rec F1 AUC Acc Prec Rec F1 AUC
CodeBERT-base 0.4965 0.4982 0.9890 0.6626 0.4421 0.6605 0.6341 0.7591 0.6909 0.7214
UniXcoder-base 0.5070 0.5036 0.9820 0.6658 0.4992 0.6745 0.6480 0.7640 0.7012 0.7290
The experimental results are shown in Tables 12 and 13. When evaluated on single-language benchmarks, models fine-tuned on the CLB dataset achieve accuracy and AUC values close to random guessing. In contrast, models fine-tuned on single-language datasets consistently obtain higher accuracy, F1 score, and AUC on both CodeBERT-base and UniXcoder-base. These results indicate that models fine-tuned on the CLB dataset have limited discriminative capability for single-language bug detection. This performance gap further suggests that features learned from cross-language interactions do not transfer well to bug detection within a single programming language.
Moreover, the performance degradation on single-language benchmarks indicates that CLB fine-tuning does not improve model effectiveness beyond its intended task. Instead, fine-tuning on the CLB dataset mainly guides CodeLMs to learn CLB features, which are largely unrelated to the characteristics of single-language bugs. Overall, this study emphasizes the necessity of constructing dedicated CLB datasets, which enable CodeLMs to learn CLB characteristics and interaction patterns essential for effective CLB detection.
@@
For CodeLMs that showed notable performance improvements, code comments provided additional context and explanations that contributed to a more accurate understanding of the code, suggesting that these CodeLMs possess relatively limited inherent code understanding capabilities. In contrast, the inference-only GPT-4o-mini baseline did not consistently benefit from the inclusion of code comments, suggesting that, for this non-fine-tuned reference setting, comment-level semantic information alone may be insufficient to substantially improve CLB detection performance without task-specific fine-tuning. Therefore, in practical applications, whether to retain code comments should be flexibly decided based on the specific task requirements. It is also advisable to consider adopting more flexible token budget strategies, such as moderately increasing the token sequence length limit or prioritizing key code content. The results from GPT-4o-mini additionally suggest that, even for relatively strong general-purpose models, input design alone may not fully compensate for the lack of task-aligned training. Nevertheless, this comparison should be interpreted cautiously, as GPT-4o-mini is used here only as a prompt-based reference baseline rather than as a representative upper bound of current frontier model performance. This finding further indicates that current CodeLMs still have considerable potential for improvement in code understanding tasks involving cross-language interactions.
@@ 5.2. Implications
Based on the study results and their analysis, we obtained a number of implications for practitioners and researchers, respectively. All implications below are derived under the experimental settings of this study, including the selected open-source CodeLMs, the constructed CLB dataset, and the fine-tuning and evaluation protocols described in Section 3.7.
@@ 5.2.1. Implications for practitioners
(1) Emphasizing the specific characteristics of CLBs. This study shows that CLBs are not only considerable in quantity but also tend to be more obscure and harder to detect compared to single-language bugs, due to the involvement of interaction mechanisms across different languages. This observation is supported by the consistently lower detection performance obtained when CodeLMs are fine-tuned only on single-language bug datasets (RQ2), indicating that CLBs exhibit characteristics that are insufficiently captured by single-language bug patterns alone. Therefore, for projects that involve cross-language interactions similar to those represented in our CLB dataset, it is advisable to explicitly consider risks related to cross-language interaction during development and testing, such as data type mismatches, inconsistent exception handling, and cross-language resource management.
(2) Paying attention to limitations of CodeLMs in CLB detection. Although CodeLMs have made significant progress in tasks such as code understanding and generation, their performance on CLB detection remains limited. As evidenced by the experimental results in RQ1 and RQ2, the detection performance of the evaluated fine-tuned CodeLMs is highly dependent on the availability of CLB-specific training data and does not consistently generalize from single-language bug datasets. Practitioners should not regard the outputs of CodeLMs as final conclusions when detecting CLBs. Instead, a multifaceted and multi-perspective detection process should be adopted by combining static analysis tools, rule-based detection methods, and manual review to comprehensively assess the possibility of CLBs. Over-reliance on the output of a single CodeLM should be avoided to enhance overall reliability of CLB detection.
(3) Considering the impact of code comments under fixed token length constraints. In this study, the results show that including code comments generally improves recall and F1 score for most CodeLMs. However, all experiments were conducted under a fixed maximum token sequence length (up to 512 tokens).
Our results therefore only indicate that, under fixed-length input constraints, the inclusion of comments provides additional contextual information but may also lead to information loss when input context is truncated. In such cases, the model is forced to make predictions based on incomplete input, which can negatively affect detection accuracy. Future work will explore the performance of CodeLMs that support longer context lengths under more powerful computational resources, in order to better capture complex cross-language interactions without aggressive truncation.
(4) Selecting an appropriate CodeLM based on resource availability and projects under detection. Under the specific experimental conditions of this study, several small CodeLMs (e.g., UniXcoder-base and GraphCodeBERT-base) achieved better or comparable performance to the evaluated larger models. It suggests that the potential advantages of large models may not be fully realized under constrained input lengths and limited fine-tuning data, which are particularly challenging in cross-language scenarios involving additional boilerplate and interaction code. Accordingly, within similar resource and input constraints, smaller CodeLMs may represent a cost-effective choice, while the effectiveness of larger CodeLMs with extended context lengths remains an open question for future investigation.
@@ 5.2.2. Implications for researchers
(1) Expanding the coverage of CodeLMs and tasks. The current study evaluates a limited number of CodeLMs, with different model architectures and parameter sizes. Future work should systematically expand the research scope to include a broader range of CodeLM architectures and large CodeLMs. Additionally, the transferability and robustness of CodeLMs across diverse cross-language scenarios should be explored to establish more comprehensive and accurate evaluation standards. Future work should also evaluate the CLB detection performance of frontier models on our dataset under zero-shot or few-shot settings.
(2) Building larger and more diverse CLB dataset. This study provides evidence of the importance of dataset quality and fine-tuning strategies. However, existing CLB datasets are still limited in terms of size, language diversity, and interaction patterns. Future work should focus on constructing larger and more comprehensive benchmark datasets that cover a wider range of PL combinations and cross-language interaction modes, in order to systematically evaluate the detection capabilities of CodeLMs of different scales and structures.
(3) Investigating the fundamental characteristics and classification of CLBs. Currently, the understanding of CLBs remains relatively shallow, and there is a lack of systematic classification standards regarding the various types of CLBs. Future research could aim to establish a structured classification system for CLBs based on their sources (e.g., memory management, exception handling, data format conversion) and manifestations (e.g., data loss, type errors, resource leaks). Such a classification system would not only facilitate more targeted CodeLM training but also support the development of finer-grained CLB detection strategies.
(4) Exploring fine-tuning strategies and data augmentation methods for CLB detection. While it is shown that increasing the volume of training data has improved CodeLM performance, efficiently leveraging limited data remains a significant challenge. Future directions include exploring self-supervised learning, domain adaptation, and cross-language contrastive learning to enhance model performance under low-resource conditions. Additionally, designing smarter data augmentation methods, such as automatically generating diverse CLB cases, could further enrich the training dataset and ultimately improve the CodeLMs’ performance in detecting CLBs.
(5) Exploring the integration of static analysis and deep learning models for CLB detection. While this study focuses on data-driven models, traditional static analysis remains a crucial approach that struggles with language boundaries. Inspired by recent neural and symbolic frameworks such as IRIS (Li et al., 2025b), future research may explore integrating deep learning with static analysis for CLB detection. By combining the rigor of traditional static analysis with the semantic understanding of deep learning models across languages, this hybrid approach presents a highly promising direction for developing advanced CLB and vulnerability detection systems.
@@ 6.2. External Validity
To answer RQ1 and RQ4, we fine-tuned CodeLMs on datasets with and without code comments, respectively, to investigate their performance in CLB detection tasks and to analyze the impact of code comments on CodeLMs effectiveness. A major threat to external validity in this setting lies in the limited parameter sizes of the CodeLMs used. In this study, we fine-tuned only 13 relatively small CodeLMs (no larger than 7B). As a result, the findings may not generalize to the performance of large models in CLB detection. Furthermore, due to computational constraints, we employed LoRA, a parameter-efficient fine-tuning (PEFT) technique, for CodeLM variants with more than 770M parameters. This may affect the generalizability of our study findings. In addition, GPT-4o-mini was included in RQ1 and RQ4 merely as an inference-only reference baseline for prompt-based, non-fine-tuned performance. Therefore, its results should not be interpreted as representing the upper bound of current frontier LLM performance, and the corresponding comparisons should be generalized with caution.
@@ 6.3. Internal Validity
A key internal threat lies in the manual analysis of CLB symptoms and root causes may introduce internal validity threats due to reliance on human judgment. To mitigate this issue, we followed an established taxonomy from prior work with minor adaptations to better fit the characteristics of the CLB dataset. Moreover, all sampled cases were independently analyzed by the first and second authors, and disagreements were resolved through discussion to reach consensus, thereby reducing subjectivity and improving internal validity.
Another important internal threat concerns the choice of hyper-parameters during the fine-tuning process. The hyper-parameters of these 13 CodeLMs during the fine-tuning process play a crucial role in the final performance on the CLB detection task. Changes in hyper-parameters can lead to different results, potentially affecting the reliability of our experimental findings. To mitigate this threat, we conducted multiple test experiments to determine the optimal hyper-parameters.
A further internal threat stems from the different fine-tuning strategies used for CodeLMs of different sizes: small CodeLMs were fine-tuned using full training, while large CodeLMs were fine-tuned using PEFT. This inconsistency may affect the fairness of model comparisons. To alleviate this threat, we explicitly reported the fine-tuning strategy used for each model in our results, and we highlighted these differences in our comparative analysis. When interpreting performance differences, we took the fine-tuning method into account to avoid drawing misleading conclusions based on absolute scores.
@@ 6.4. Reliability
To further ensure the reliability and reproducibility of our experiments, we publicly shared the data collection standards and the final CLB dataset (Li et al., 2025a). Detailed descriptions were provided for each step of the data collection process (see Section 3.3), including repository selection criteria (e.g., repositories with more than 500 stars, cross-language collaboration, and relevant bug issues) and the procedure of data processing. These steps ensured the standardization of data collection, and all relevant parameters (such as the repository star threshold and the minimum proportion of code in each PL) were clearly specified and justified in the methodology.
In the dataset partitioning phase, we used a fixed random seed to ensure consistent splitting of the training, validation, and test sets, thereby eliminating randomness that could affect the results. Additionally, we made our experimental scripts publicly available(Li et al., 2025a) to guarantee the reproducibility of our experiments. Thus, with these measures, the threats to reliability are alleviated.
@@ 7. Conclusions and Future Work
In this work, we designed CLCFinder, a tool for identifying cross-language code, and constructed a real-world dataset for CLB detection from GitHub. Based on this dataset, we conducted experiments to evaluate the performance of widely-used CodeLMs in the CLB detection task. The results show significant performance variation across different CodeLMs, with small CodeLMs (the parameter size is 220M or less) tending to achieve better performance than large CodeLMs within our experimental setting. To further validate the differences between our CLB dataset and single-language bug datasets, we fine-tuned CodeLMs on SOTA single-language bug detection datasets and applied these fine-tuned CodeLMs to CLB detection. The results demonstrate significant differences in feature characteristics between single-language bug and CLB datasets, validating the effectiveness and necessity of our CLB dataset. However, although the mainstream CodeLMs have demonstrated preliminary effectiveness in CLB detection, they are still primarily used in research settings and are not yet ready for real-world applications. We then explored factors affecting the performance of CodeLMs in CLB detection. The results show that increasing dataset size improves CodeLM performance, while the impact of token sequence length varies across CodeLMs. Some CodeLMs perform better with a shorter token sequence length, while others perform better with a longer token sequence length. Finally, we investigated the impact of code comments on model performance in CLB detection. Experimental results indicate that code comments affect CodeLMs differently - some CodeLMs perform better with commented code, some CodeLMs show no significant performance change, and others perform worse.
@@
For future work, we will advance research on CLB detection along several key directions. First, we intend to conduct a more in-depth analysis of the CLB dataset itself. This includes investigating the characteristics of CLBs, such as common interaction patterns across language boundaries, as well as analyzing CLB fix characteristics to better understand how CLBs are resolved in practice. Such analyses will provide deeper insights into the nature of CLBs and help inform the design of more effective detection and repair approaches. Second, we plan to expand our dataset beyond Python and Java to include more PLs. This will enable more comprehensive CLB studies and improve the generalizability of detection models. Additionally, since our dataset includes bugs related to high-severity security vulnerabilities, we aim to explore cross-language security vulnerability detection in future research. Third, while this study focused on relatively small CodeLMs (no larger than 7B), we plan to investigate the performance of large CodeLMs in CLB detection. We also intend to conduct more in-depth analysis to better understand the factors that influence CLB detection performance, such as model architecture and language-specific features. Finally, we will examine the effect of longer input sequences and larger datasets on model performance. In this study, input token sequence length was limited to 512, which may have restricted the CodeLM’s ability to capture long-range dependencies. We plan to expand our dataset in order to study the impact of dataset size on CLB detection.
@@
The replication package of this work has been made available at (Li et al., 2025a).